OpenStudy (anonymous):

please help on this one: find the sum of this rational expression x-1/x-2 + x-6/x^2-4

OpenStudy (anonymous):

f (x) = (x - 1)/(x - 2) + (x - 6)/(x^2 - 4) Common denominator: (x^2 - 4) = (x -2)(x + 2) Numerator of f(x): (x -1)(x + 2) + x - 6 = x^2 + 2x - 8 = (x - 2)(x + 4) Denominator of f(x): (x - 2)(x + 2). After simplification, f(x) = ......